# Mavors: Multi-granularity Video Representation for Multimodal Large Language Model

### v1

@@ 5.3 Ablation Studies
We conduct a series of ablation studies to validate our model design. Given the extensive training time required for the full training paradigm, these ablations utilize standard compositive datasets and train various versions up to the completion of Stage 2. Specifically, Stage 1 employs LLaVA-Pretrain-558K [53] and LLaVA-Hound-Pretrain [122]; Stage 1.5 uses M4-Instruct [44] and ShareGPT4o [16]; and Stage 2 utilizes LLaVA-OneVision and LLaVA-Video. This approach reduces the duration of a full training cycle to under 24 hours with 64 A800 GPUs. Performance is subsequently monitored using MMMU, MathVista, and CapsBench for image understanding capabilities, and Video-MME, Vinoground, and DREAM-1K for video understanding capabilities.
Effect of the Number of Frames in a Video Chunk. We conduct experiments with four settings, varying a parameter $F$ with values of 4, 8, 16, and 32. Upon the preliminary study evaluating video captioning performance on the validation set of KVQ [63], we observe that configurations with $F=8$ or $F=16$ yield more accurate and comprehensive captions. To ensure exposure to richer visual information, we finalize the $F=16$ setting. We further evaluate these four model variants on six benchmark datasets in Figure 9. On image-based tasks, we observe a marginal improvement in performance metrics with increasing $F$. We hypothesize that this improvement stems from the model’s increased exposure to individual frames during video processing when $F$ is larger, thereby enhancing its image understanding capabilities. Conversely, for video understanding tasks, performance degrades significantly for $F=4$ due to insufficient temporal information and for $F=32$, likely due to excessive information compression.
@@ 6 Conclusion
Image Caption LAION (EN 6.7M, ZH 3.2M) [81], Conceptual Captions (7.3M) [83], SBU (0.8M) [73], COYO (11M) [7], WuKong (2.9M) [25], LAION COCO (16M) [1], OMEGA Image Caption (79M) [39]
Video Caption InternVid-10M-FLT (1.6M) [102], Panda-70M (0.9M) [12], OMEGA Video Caption (4M) [39]
@@
Image Caption Met-meme [107], PD12M [68], dalle3 [71], GBC10M [30], DenseFusion-1M [48], GameBunny [89], MERMAID [92], CC12M (1M) [9], BLIP3 [3], AllSeeingV2 [100]
Video Caption ChronoMagic [113], VideoChatGPT [67], YouCook2 [127], CelebV [111], SthSthV2 [24], MiraData [35], Hacs [125], OpenVid-1M [72], Kinetics_700 [8], ShareGPT4Video [11], Vript [109], Shot2Story [27], ShareGemini [82]
Question Answering MMDU [59], MMiT [70]
Knowledge Wikipedia [21], Wikimedia [21], WIT [87]
Code WebSight [42]
OCR LSVT [88], ArT [15], DocMatix [41]
Interleaved OBELICS [40], PIN [97]
Mixed-Task Dataset MMInstruct [57], LVD-2M [106], MMEvol [64]
@@
Instruction Countix [18], VideoChat [45], Videogpt+ [66], Openmathinstruct-2 (2M) [93], RepCountA [31], Vidgen-1m [91], CompCap [13], Metamath [112], Llava-Onevision [43], Anytext (0.3M) [94], Llava-Video [124], S-MiT [69], LSMDC [80], Infinity-MM [26], Mantis [33], ShareGPT4V [10], CinePile [79], LLaVA-Hound [122]
Grounding GRIT [75], RefCOCO [36]
Temporal Grounding GroundedVideoLLM [95]
@@
Open-ended QA Llava-Video [124] (10K)
Image Caption Llava-Onevision [43] (10K), DenseFusion-1M [48] (10K)
Video Caption WebVid [5] (8K), Kinetics_700 [8] (8K), OOPS [19] (4K)
@@ A Training Datasets
We have also curated two datasets from the OMEGA project [39], the OMEGA Image Caption (containing 79M samples) and OMEGA Video Caption (containing 4M samples), by sampling videos and images along with their corresponding titles and captions. These two datasets are utilized in the first stage of our model training.
For certain datasets that either lack captions or only possess low-quality ones, for example, CC12M [9], CelebV [111], Hacs [125], and Kinetics_700 [8], we carefully designed a pipeline to generate high-quality captions. Initially, we utilized Qwen2VL-72B [98], InternVL2.5-78B-MPO [14] and Tarsier-34B [96] (video only) to describe these samples in detail. Subsequently, we used DeepSeek-R1-Distill-Llama-70B [17] to amalgamate captions generated by different models while attempting to resolve all inconsistencies using its COT capabilities. The captions produced by this process generally demonstrated superior quality and comprehensibility.
We observed that many composite datasets incorporate content from established standalone datasets, leading to potential data redundancy. To address this, we implemented a deduplication process for identical samples (images or videos). Specifically, we calculated the Perplexity (PPL) of the associated text using the Qwen2VL-72B [98] model, distinguishing between QA and Captioning tasks. For duplicate visual content within QA tasks, we retained the two samples exhibiting the lowest text PPL scores. For Captioning tasks, one sample was randomly selected from the two with the lowest PPL for inclusion in our training set.
@@ B Analysis on the Inference Costs
We evaluate the inference performance of Qwen2.5VL-7B and Mavors-7B using an NVIDIA GeForce RTX 4090 GPU. Initially, we measure the execution time of the model.generate function via the standard HuggingFace implementation (with FlashAttention-2 enabled) to capture the core model execution time, excluding video preprocessing. Table 7 summarizes the inference times for both models on the DREAM-1K and CapsBench video captioning tasks. The results show that Mavors’ more efficient video representation reduces both the ViT computations and the language model’s context window requirements. Consequently, Mavors-7B demonstrates significant speed improvements on video understanding tasks, achieving 2.7x faster prefill and 1.4x faster decoding compared to Qwen2.5VL-7B. Furthermore, integrating the vLLM inference framework with overlapping vision preprocessing enables 2.5s per image in CapsBench and 3.7s per video in DREAK-1K, reducing from about 13s per image and 20s per video respectively. These findings indicate that Mavors provides an economical solution for scenarios requiring frequent or high-volume multimodal model inference.
@@ C Details of Experiments
Evaluation Setup. To ensure a standardized and reproducible evaluation, we conduct experiments on both open-source and closed-source models using consistent protocols. For open-source models, we adopt the lmms-eval framework [119], which offers a unified pipeline tailored for benchmarking MLLMs. All open-source models are evaluated using the officially released checkpoints to preserve the integrity of reported results. To maintain experimental stability, we fix the decoding strategy to greedy decoding, set the maximum number of generated tokens to 1024. Image and video resolution, along with other preprocessing settings, follow the default configurations provided by the lmms-evak framework or the respective model implementations. For closed-source models, including Gemini-1.5-Pro-002 [23] and GPT-4o-20240806 [32], we access them through their official APIs. Due to the restricted controllability over decoding parameters, we adopt the default generation settings provided by each platform. For benchmarks requiring GPT-based automatic scoring, such as those involving instruction-following or open-ended generation tasks, we follow the evaluation protocol described in the original benchmark papers or apply the default settings specified by the lmms-eval framework to select the judge model. Specifically, for MathVista [62], we use GPT-4-Turbo (1106) as the judge model. For CapsBench [52] and MMMU [114], we adopt GPT-4o (20240806), while for DREAM-1K [96], we follow the original benchmark and employ GPT-3.5-Turbo (0125) to perform automatic scoring. These choices align with the evaluation protocols used in the respective benchmark papers, ensuring fair and comparable results across models.
@@
For image benchmarks, we compare against two leading proprietary models, GPT-4o [32] and Gemini-1.5-Pro [23]. GPT-4o, developed by OpenAI, is capable of processing text, images, and audio in a unified manner and has demonstrated strong performance in visual reasoning tasks. Gemini, developed by Google DeepMind, similarly integrates multimodal capabilities and excels in scenarios requiring complex cross-modal understanding. We also include a range of high-performing open-source MLLMs in our comparison. These include CogVLM2 [29], a model optimized for visual-language understanding in dynamic contexts; GLM-4V [29], which extends the GLM architecture with strong visual recognition capabilities; LLaVA-OneVision [43], a widely recognized open-source MLLM that integrates a collection of high-quality multimodal datasets, advanced training strategies, and refined model designs to achieve strong performance across image-based benchmarks; InternVL2.5 [14], which is an advanced MLLM series developed by Shanghai Artificial Intelligence Laboratory. Building upon the architecture of InternVL2, it introduces significant enhancements in training strategies and data quality; DeepSeek-VL2 [105], an MoE-based model balancing scalability and accuracy; and Qwen2.5-VL [4], a model that significantly enhance general image recognition capabilities, expanding to a vast array of categories, including plants, animals, landmarks, and various products. It also excels in precise object localization, advanced text recognition, and document parsing.
@@
For video benchmarks, we select four representative categories of baseline models, each exemplifying distinct video processing strategies. The first category includes models that employ sparse frame sampling with high performance, such as NVILA [61] and LLaVA-Video [123], which focus on selecting key frames to reduce computational overhead while maintaining contextual understanding. NVILA, developed by NVIDIA, utilizes a “scale-then-compress” paradigm that first increases spatial and temporal resolutions and then compresses visual tokens, enabling efficient processing of high-resolution images and long videos. LLaVA-Video improves video understanding by introducing a high-quality synthetic dataset, LLaVA-Video-178K [123], specifically designed for video instruction-following tasks. Models like Qwen2.5-VL [4] and Oryx-1.5 [60] adopt dense frame sampling at lower resolutions to achieve a trade-off between information richness and efficiency (we set at most 768 frames in our experiments). Oryx-1.5 is a unified MLLM designed to flexibly and efficiently handle visual inputs with varying spatial scales and temporal lengths, making it well-suited for processing both high-resolution images and extended video sequences. In addition, we include models such as VideoChat-Flash [47] and VideoLLaMA3 [116], which apply dense sampling combined with token compression to handle long video sequences effectively (up to 1000 frames in our experiments). VideoChat-Flash leverages this strategy to mitigate the computational overhead introduced by dense sampling, enabling effective handling of long-duration inputs without sacrificing performance. Similarly, VideoLLaMA3 integrates token compression with dense sampling to reduce input redundancy, thereby enhancing the model’s ability to understand and reason over extended video content. Finally, we include Slow-fast MLLM [84], which employs a specialized dual-pathway sampling mechanism to capture temporal dynamics at multiple granularities. By processing visual inputs through both slow and fast pathways, the model effectively models temporal variations across different timescales.
@@
We adopt MMMU [114], MathVista [62], AI2D [37], and CapsBench [52] as representative image benchmarks, covering a broad range of visual understanding and reasoning tasks.
@@
We conduct evaluations on a diverse set of video benchmarks, including MMWorld [28], PerceptionTest [74], Video-MME [22], MLVU [126], MVBench [46], EventHallusion [117], TempCompass [58], VinoGround [118], and DREAM-1K [96].
@@ D Needle in a Haystack Test
Inspired by the design in LongVA [121], we conduct Needle-in-a-Haystack (NIAH) test. We adopt a chunk-level NIAH evaluation scheme, which primarily focuses on evaluating the model’s comprehension accuracy when a target frame is inserted into different video chunks. We utilize 10 short-duration and 10 medium-duration videos from the Video-MME benchmark. We examine the model’s performance across video lengths ranging from 1 to 60 chunks. Recall that 60 chunks correspond to 960 frames. For a given number of chunks $c_{\text{V}}$, we performed 50*$c_{\text{V}}$ trials. In each trial, we randomly select a video, an image (the ‘needle’), a specific chunk within the video, and a frame position within that chunk. The selected image then replaces the original frame at the chosen position. Notably, after selecting a video, we first apply accelerating playback (temporal subsampling) to ensure the video frames precisely fit into $c_{\text{V}}$ chunks. Figure 11 illustrates the accuracy results. As observed, perfect accuracy is achieved within the model’s training window length (32 chunks). Moreover, Mavors maintains strong accuracy even as the number of chunks increases beyond this window. This experiment indicates that Mavors can provide reliable understanding for videos of reasonable duration, provided essential frame information is not lost during the accelerating playback process.

### v2

@@ 5.3 Ablation Studies
We conduct a series of ablation studies to validate our model design. Given the extensive training time required for the full training paradigm, these ablations utilize standard compositive datasets and train various versions up to the completion of Stage 2. Specifically, Stage 1 employs LLaVA-Pretrain-558K [liu2023llava15] and LLaVA-Hound-Pretrain [llava-hound]; Stage 1.5 uses M4-Instruct [li2024llavanextinterleave] and ShareGPT4o [cui2025comprehensive]; and Stage 2 utilizes LLaVA-OneVision and LLaVA-Video. This approach reduces the duration of a full training cycle to under 24 hours with 64 GPUs. Performance is subsequently monitored using MMMU, MathVista, and CapsBench for image understanding capabilities, and Video-MME, Vinoground, and DREAM-1K for video understanding capabilities.
Effect of the Number of Frames in a Video Chunk. We conduct experiments with four settings, varying a parameter $F$ with values of 4, 8, 16, and 32. Upon the preliminary study evaluating video captioning performance on the validation set of KVQ [lu2024kvq], we observe that configurations with $F=8$ or $F=16$ yield more accurate and comprehensive captions. To ensure exposure to richer visual information, we finalize the $F=16$ setting. We further evaluate these four model variants on six benchmark datasets in Figure 9. On image-based tasks, we observe a marginal improvement in performance metrics with increasing $F$. We hypothesize that this improvement stems from the model’s increased exposure to individual frames during video processing when $F$ is larger, thereby enhancing its image understanding capabilities. Conversely, for video understanding tasks, performance degrades significantly for $F=4$ due to insufficient temporal information and for $F=32$, likely due to excessive information compression.
@@ 6 Conclusion
Image Caption LAION (EN 6.7M, ZH 3.2M) [laion400m], Conceptual Captions (7.3M) [sharma2018conceptual], SBU (0.8M) [ordonez2011im2text], COYO (11M) [coyo-700m], WuKong (2.9M) [gu2022wukong], LAION COCO (16M) [lacion-coco], OMEGA Image Caption (79M) [omegaproject]
Video Caption InternVid-10M-FLT (1.6M) [wang2023internvid], Panda-70M (0.9M) [chen2024panda], OMEGA Video Caption (4M) [omegaproject]
@@
Image Caption Met-meme [xu2022metmeme], PD12M [meyer2024public], dalle3 [dalle3], GBC10M [hsieh2024graph], DenseFusion-1M [li2024densefusion], GameBunny [taesiri2024videogamebunny], MERMAID [toh2023mermaid], CC12M (1M) [changpinyo2021cc12m], BLIP3 [awadalla2024blip3], AllSeeingV2 [wang2024all]
Video Caption ChronoMagic [yuan2024chronomagic], VideoChatGPT [Maaz2023VideoChatGPT], YouCook2 [zhou2018towards], CelebV [yu2023celebv], SthSthV2 [goyal2017something], MiraData [ju2024miradata], Hacs [zhao2019hacs], OpenVid-1M [nan2024openvid], Kinetics_700 [carreira2019short], ShareGPT4Video [chen2024sharegpt4video], Vript [yang2024vript], Shot2Story [han2023shot2story20k], ShareGemini [sharegemini]
Question Answering MMDU [liu2024mmdu], MMiT [monfort2021multi]
Knowledge Wikipedia [wikidump], Wikimedia [wikidump], WIT [srinivasan2021wit]
Code WebSight [laurençon2024unlocking]
OCR LSVT [sun2019icdar], ArT [chng2019icdar2019], DocMatix [laurenccon2024building]
Interleaved OBELICS [laurencon2023obelics], PIN [wang2024pin]
Mixed-Task Dataset MMInstruct [liu2024mminstruct], LVD-2M [xiong2024lvd], MMEvol [luo2024mmevol]
@@
Instruction Countix [Dwibedi_2020_CVPR], VideoChat [2023videochat], Videogpt+ [maaz2024videogpt+], Openmathinstruct-2 (2M) [toshniwal2024openmathinstruct], RepCountA [hu2022transrac], Vidgen-1m [tan2024vidgen], CompCap [chen2024compcap], Metamath [yu2023metamath], Llava-Onevision [li2024llava], Anytext (0.3M) [tuo2023anytext], Llava-Video [zhang2024video], S-MiT [monfort2021spoken], LSMDC [rohrbach2017movie], Infinity-MM [gu2024infinity], Mantis [Jiang2024MANTISIM], ShareGPT4V [chen2023sharegpt4v], CinePile [rawal2024cinepile], LLaVA-Hound [llava-hound]
Grounding GRIT [Kosmos2], RefCOCO [kazemzadeh2014referitgame]
Temporal Grounding GroundedVideoLLM [wang2024grounded]
@@
Open-ended QA Llava-Video [zhang2024video] (10K)
Image Caption Llava-Onevision [li2024llava] (10K), DenseFusion-1M [li2024densefusion] (10K)
Video Caption WebVid [bain2021frozen] (8K), Kinetics_700 [carreira2019short] (8K), OOPS [epstein2019oops] (4K)
@@ A Training Datasets
We have also curated two datasets from the OMEGA project [omegaproject], the OMEGA Image Caption (containing 79M samples) and OMEGA Video Caption (containing 4M samples), by sampling videos and images along with their corresponding titles and captions. These two datasets are utilized in the first stage of our model training.
For certain datasets that either lack captions or only possess low-quality ones, for example, CC12M [changpinyo2021cc12m], CelebV [yu2023celebv], Hacs [zhao2019hacs], and Kinetics_700 [carreira2019short], we carefully designed a pipeline to generate high-quality captions. Initially, we utilized Qwen2VL-72B [Qwen2VL], InternVL2.5-78B-MPO [chen2024expanding] and Tarsier-34B [wang2024tarsier] (video only) to describe these samples in detail. Subsequently, we used DeepSeek-R1-Distill-Llama-70B [deepseekai2025deepseekr1] to amalgamate captions generated by different models while attempting to resolve all inconsistencies using its COT capabilities. The captions produced by this process generally demonstrated superior quality and comprehensibility.
We observed that many composite datasets incorporate content from established standalone datasets, leading to potential data redundancy. To address this, we implemented a deduplication process for identical samples (images or videos). Specifically, we calculated the Perplexity (PPL) of the associated text using the Qwen2VL-72B [Qwen2VL] model, distinguishing between QA and Captioning tasks. For duplicate visual content within QA tasks, we retained the two samples exhibiting the lowest text PPL scores. For Captioning tasks, one sample was randomly selected from the two with the lowest PPL for inclusion in our training set.
@@ B Analysis on the Inference Costs
We evaluate the inference performance of Qwen2.5VL-7B and Mavors-7B using an GPU. Initially, we measure the execution time of the model.generate function via the standard HuggingFace implementation (with FlashAttention-2 enabled) to capture the core model execution time, excluding video preprocessing. Table 7 summarizes the inference times for both models on the DREAM-1K and CapsBench video captioning tasks. The results show that Mavors’ more efficient video representation reduces both the ViT computations and the language model’s context window requirements. Consequently, Mavors-7B demonstrates significant speed improvements on video understanding tasks, achieving 2.7x faster prefill and 1.4x faster decoding compared to Qwen2.5VL-7B. Furthermore, integrating the vLLM inference framework with overlapping vision preprocessing enables 2.5s per image in CapsBench and 3.7s per video in DREAK-1K, reducing from about 13s per image and 20s per video respectively. These findings indicate that Mavors provides an economical solution for scenarios requiring frequent or high-volume multimodal model inference.
@@ C Details of Experiments
Evaluation Setup. To ensure a standardized and reproducible evaluation, we conduct experiments on both open-source and closed-source models using consistent protocols. For open-source models, we adopt the lmms-eval framework [zhang2024lmms], which offers a unified pipeline tailored for benchmarking MLLMs. All open-source models are evaluated using the officially released checkpoints to preserve the integrity of reported results. To maintain experimental stability, we fix the decoding strategy to greedy decoding, set the maximum number of generated tokens to 1024. Image and video resolution, along with other preprocessing settings, follow the default configurations provided by the lmms-evak framework or the respective model implementations. For closed-source models, including Gemini-1.5-Pro-002 [gemini] and GPT-4o-20240806 [hurst2024gpt], we access them through their official APIs. Due to the restricted controllability over decoding parameters, we adopt the default generation settings provided by each platform. For benchmarks requiring GPT-based automatic scoring, such as those involving instruction-following or open-ended generation tasks, we follow the evaluation protocol described in the original benchmark papers or apply the default settings specified by the lmms-eval framework to select the judge model. Specifically, for MathVista [mathvista], we use GPT-4-Turbo (1106) as the judge model. For CapsBench [liu2024playground] and MMMU [yue2023mmmu], we adopt GPT-4o (20240806), while for DREAM-1K [wang2024tarsier], we follow the original benchmark and employ GPT-3.5-Turbo (0125) to perform automatic scoring. These choices align with the evaluation protocols used in the respective benchmark papers, ensuring fair and comparable results across models.
@@
For image benchmarks, we compare against two leading proprietary models, GPT-4o [hurst2024gpt] and Gemini-1.5-Pro [gemini]. GPT-4o, developed by OpenAI, is capable of processing text, images, and audio in a unified manner and has demonstrated strong performance in visual reasoning tasks. Gemini, developed by Google DeepMind, similarly integrates multimodal capabilities and excels in scenarios requiring complex cross-modal understanding. We also include a range of high-performing open-source MLLMs in our comparison. These include CogVLM2 [hong2024cogvlm2], a model optimized for visual-language understanding in dynamic contexts; GLM-4V [hong2024cogvlm2], which extends the GLM architecture with strong visual recognition capabilities; LLaVA-OneVision [li2024llava], a widely recognized open-source MLLM that integrates a collection of high-quality multimodal datasets, advanced training strategies, and refined model designs to achieve strong performance across image-based benchmarks; InternVL2.5 [chen2024expanding], which is an advanced MLLM series developed by Shanghai Artificial Intelligence Laboratory. Building upon the architecture of InternVL2, it introduces significant enhancements in training strategies and data quality; DeepSeek-VL2 [deepseek_vl2], an MoE-based model balancing scalability and accuracy; and Qwen2.5-VL [qwen2.5vl], a model that significantly enhance general image recognition capabilities, expanding to a vast array of categories, including plants, animals, landmarks, and various products. It also excels in precise object localization, advanced text recognition, and document parsing.
@@
For video benchmarks, we select four representative categories of baseline models, each exemplifying distinct video processing strategies. The first category includes models that employ sparse frame sampling with high performance, such as NVILA [liu2024nvila] and LLaVA-Video [llava-video-sft], which focus on selecting key frames to reduce computational overhead while maintaining contextual understanding. NVILA, developed by NVIDIA, utilizes a “scale-then-compress” paradigm that first increases spatial and temporal resolutions and then compresses visual tokens, enabling efficient processing of high-resolution images and long videos. LLaVA-Video improves video understanding by introducing a high-quality synthetic dataset, LLaVA-Video-178K [llava-video-sft], specifically designed for video instruction-following tasks. Models like Qwen2.5-VL [qwen2.5vl] and Oryx-1.5 [liu2024oryx] adopt dense frame sampling at lower resolutions to achieve a trade-off between information richness and efficiency (we set at most 768 frames in our experiments). Oryx-1.5 is a unified MLLM designed to flexibly and efficiently handle visual inputs with varying spatial scales and temporal lengths, making it well-suited for processing both high-resolution images and extended video sequences. In addition, we include models such as VideoChat-Flash [videochat-flash] and VideoLLaMA3 [zhang2025videollama3], which apply dense sampling combined with token compression to handle long video sequences effectively (up to 1000 frames in our experiments). VideoChat-Flash leverages this strategy to mitigate the computational overhead introduced by dense sampling, enabling effective handling of long-duration inputs without sacrificing performance. Similarly, VideoLLaMA3 integrates token compression with dense sampling to reduce input redundancy, thereby enhancing the model’s ability to understand and reason over extended video content. Finally, we include Slow-fast MLLM [shi2025slow], which employs a specialized dual-pathway sampling mechanism to capture temporal dynamics at multiple granularities. By processing visual inputs through both slow and fast pathways, the model effectively models temporal variations across different timescales.
@@
We adopt MMMU [yue2023mmmu], MathVista [mathvista], AI2D [kembhavi2016diagram], and CapsBench [liu2024playground] as representative image benchmarks, covering a broad range of visual understanding and reasoning tasks.
@@
We conduct evaluations on a diverse set of video benchmarks, including MMWorld [hemmworld], PerceptionTest [patraucean2023perception], Video-MME [videomme], MLVU [MLVU], MVBench [li2024mvbench], EventHallusion [zhang2024eventhallusion], TempCompass [tempcompass], VinoGround [vinoground], and DREAM-1K [wang2024tarsier].
@@ D Needle in a Haystack Test
[…]